OpenStudy (anonymous):

Simplify: the square root of 2 (x − the square root of 5) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

able to get anywhere?

OpenStudy (anonymous):

No, I'm not sure where to start first.

jimthompson5910 (jim_thompson5910):

the first step would be to distribute the \(\large \sqrt{2}\) term through

jimthompson5910 (jim_thompson5910):

what do you get when you do that

OpenStudy (anonymous):

2x and .. i'm not sure if i can combine a radical w/ a normal number..

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{2}(x - \sqrt{5})\] \[\large \sqrt{2}*x - \sqrt{2}*\sqrt{5}\] \[\large x*\sqrt{2} - \sqrt{2}*\sqrt{5}\] what's next?

OpenStudy (cggurumanjunath):